### v1

# Electric and spin current vortices in altermagnets

Abstract
Altermagnets constitute a class of collinear magnets with momentum-dependent spin splitting and vanishing net magnetization. Direct observation of the characteristic altermagnetic spin splitting, however, remains challenging. Indirect signatures can be obtained via transport studies, which so far have only considered homogeneous fields driving currents. We propose to leverage nonuniform electric fields and spin density gradients to probe the shape and the spin polarization of altermagnetic Fermi surfaces via transport measurements. By using both a semiclassical Boltzmann approach and a lattice Keldysh formalism, we show that altermagnets excite swirling electric and spin currents whose profiles depend on the relative orientation of altermagnetic lobes with respect to the sample boundaries. These currents can be measured via magnetometry techniques. Unlike previous proposals considering the hydrodynamic regime of transport, swirling currents are observed even in the Ohmic regime and rely exclusively on the altermagnetic spin splitting, with no swirls observed in ferromagnets. The electric and spin current vortices predicted here provide a unique altermagnetic signature in an experimentally accessible setup.
Introduction.— Transport of charge and spin represents one of the fundamental types of measurements in condensed matter physics that are indispensable in identifying and investigating new materials. Among such materials that have recently attracted much attention are magnetic materials with anisotropic nonrelativistically spin-split energy bands [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Known as altermagnets [2, 8], such materials are characterized by a combined symmetry including lattice rotations and spin reversal. Altermagnets preserve the inversion symmetry but break the time-reversal symmetry (TRS).
Altermagnetism was predicted in several materials such as MnTe, a few atom layers of Ru2O, Mn5Si3, MnF2, V2Se2O, V2Te2O, CrO, CrSb, and intercalated transition metal dichalcogenides CoNb4Se8 [8, 11]. Signatures of spin-split energy bands were observed in angle-resolved photoemission spectroscopy (ARPES) [12, 13, 14, 15, 16, 17, 18, 19, 20, 21] and circular dichroism [22]. However, ARPES measurements generically suffer from low resolution and strong reliance on the comparison between the experimental data and the ab initio calculations.
Transport experiments in altermagnets are mostly focused on the observation of the anomalous Hall effect (AHE) [23, 24, 25]. The AHE, however, may originate from different sources, such as a reduced symmetry at the interface [26], domain walls [27], or spin canting [28, 29], see also Ref. [30]. Therefore, the AHE alone cannot be used to unambiguously prove the altermagnetic nature of a material. Thus, other types of probes are necessary to rigorously prove the realization of the altermagnetic state.
In this Letter, we propose to use point contacts in altermagnets to generate swirling electric and spin currents whose patterns are directly linked to the symmetry of the altermagnetic spin splitting and the geometry of the sample. While vortices in electron fluids are usually considered as a key indicator of electron hydrodynamics [31], swirling currents in our work appear in the Ohmic transport regime. The proposed effects can be readily investigated via available experimental methods. In particular, the predicted swirling electric currents can be probed via quantum spin magnetometry [32, 33], whereas the distribution of the electrochemical potential can be measured via scanning tunneling potentiometry [34]. Local spin currents can also be probed via techniques similar to the non-local spin valve measurement [35, 36].
Continuum model and kinetic equations.— We start with defining the low-energy continuum model of altermagnets. We focus on two-dimensional (2D) altermagnets as they provide the most straightforward way to visualize electric and spin currents; the generalization of our formalism to three-dimensional cases is straightforward.
The electron Hamiltonian reads as
$H=t_{0}k^{2}+\sigma_{z}J(\mathbf{k})-\hat{\mu}.$ (1)
Here, $\mathbf{k}$ is the momentum vector, $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$, $\sigma_{z}$ is the Pauli matrix in the spin space, $J(\mathbf{k})$ is an effective exchange field, $\hat{\mu}=\text{diag}\left(\mu_{+},\mu_{-}\right)$, where $\mu_{\lambda}$ is the spin-resolved chemical potential for the spin projection $\lambda=\pm$  [^1] The spin-dependent effective chemical potential can be achieved by applying a weak magnetic field via the Zeeman term. In 2D systems, orbital effects can be ignored if the field is in the plane of the material., and parameter $t_{0}$ corresponds to the inverse effective mass of quasiparticles. To obtain finite Fermi surfaces, we assume that $t_{0}k^{2}>\left|J(\mathbf{k})\right|$.
The dispersion relation of electron quasiparticles is
$\varepsilon_{\lambda}=t_{0}k^{2}+\lambda J(\mathbf{k}).$ (2)
The corresponding Fermi surfaces for a $d$-wave altermagnet are schematically shown by the red ($\lambda=+$) and blue ($\lambda=-$) ellipses in Fig. 1.
Figure 1: Schematic setup of a magnet with point-like contacts. The arrows illustrate the flow of electric (spin) currents induced by the applied spin density (electric potential) difference at the contacts. The upper half corresponds to a $d$-wave altermagnet and the lower half represents a ferromagnet; the corresponding Fermi surfaces with spin-up and spin-down polarized bands are shown in red and blue. Despite the system being diffusive and non-interacting, current vortices arise in the altermagnetic case, unlike the ferromagnetic one.
To address the transport properties of the continuum model, we employ the standard semiclassical approach, see Appendix for details and definitions. The Boltzmann equation leads to the standard continuity equation and the following equation for the spin-resolved current $\mathbf{j}_{\lambda}(t,\mathbf{r})$ density:
$\frac{\partial j_{\lambda,l}}{\partial t}+\frac{\sigma_{il}}{\tau}\nabla_{i}\bar{\phi}_{\lambda}=-\frac{j_{\lambda,l}}{\tau}.$ (3)
In writing this equation, we assumed weak deviations from equilibrium and introduced the conductivity tensor
$\sigma_{ij}=2e\tau\rho^{(0)}_{\lambda}\left(t_{0}\delta_{ij}+\lambda T_{\lambda,ij}\right),$ (4)
where
$T_{\lambda,ij}=\frac{e}{2\rho^{(0)}_{\lambda}}\int\frac{d^{2}k}{(2\pi)^{2}}\left[\partial_{k_{i}}\partial_{k_{j}}J(\mathbf{k})\right]f_{\lambda}^{(0)}(\mathbf{k})$ (5)
is the part of the conductivity tensor determined by the effective exchange field, $f^{(0)}_{\lambda}(\mathbf{k})=1/\left[e^{\left(\varepsilon_{\lambda}-\mu_{\lambda}\right)/T}+1\right]$ is the equilibrium distribution function, $\rho^{(0)}_{\lambda}$ is the equilibrium charge density, and $e\bar{\phi}_{\lambda}=\delta\mu_{\lambda}+e\phi$ is the effective spin-resolved electrochemical potential with $\phi$ being electric potential and $\delta\mu_{\lambda}$ being the spin-resolved deviation of the chemical potential from its equilibrium value. Elastic scattering is quantified by the scattering time $\tau$  [^2] Since the spin-orbital coupling is usually weak in altermagnets, we assume the spin-flipping scattering to be also weak in altermagnets. We checked that such processes do not qualitatively affect the shape of the current streamlines [39]..
Equation (3) is general and can be applied both for static and dynamic perturbations. In this Letter, we focus on the stationary transport regime, where Eq. (3) reduces to
$\nabla_{j}\left(\sigma_{ij}\nabla_{i}\bar{\phi}_{\lambda}\right)=0.$ (6)
This differential equation should be supplemented with boundary conditions. For example, one could fix $\bar{\phi}_{\lambda}$ at the contacts or require a constant injected current. For a known $\sigma_{ij}$, Eq. (6) is a partial differential equation that can be solved numerically or, in certain cases, such as a ribbon with the fixed values of the current at the point-like contacts, see the Supplemental Material (SM) [39], analytically.
Vorticity in the Ohmic transport regime.— Before delving into the numerical solutions to Eq. (6), let us present intuitive arguments supporting the possibility of vortices in the Ohmic transport regime. We start with the analog of the vorticity equation in fluid mechanics [40], where we define the spin-resolved analog of vorticity as $\bm{\omega}_{\lambda}=\bm{\nabla}\times\mathbf{j}_{\lambda}$. Assuming the absence of the equilibrium spin density, $\mu_{\lambda}=\mu$, the equations for the electric $\bm{\omega}_{\rm el}=\sum_{\lambda}\bm{\omega}_{\lambda}$ and spin $\bm{\omega}_{\rm sp}=\sum_{\lambda}\lambda\bm{\omega}_{\lambda}$ vorticities follow from Eq. (3)
$\frac{\partial\omega_{{\rm el},j}}{\partial t}+\frac{\omega_{{\rm el},j}}{\tau}=-2e\rho^{(0)}\epsilon_{jlm}T_{im}\nabla_{l}\nabla_{i}\bar{\phi}_{\rm sp},$ (7)
$\frac{\partial\omega_{{\rm sp},j}}{\partial t}+\frac{\omega_{{\rm sp},j}}{\tau}=-2e\rho^{(0)}\epsilon_{jlm}T_{im}\nabla_{l}\nabla_{i}\bar{\phi}_{\rm el},$ (8)
where $\epsilon_{jlm}$ is the fully antisymmetric tensor. Without the off-diagonal terms in the conductivity tensor $T_{\lambda,i\neq m}$, vorticity is trivial. Such terms can originate, e.g., from the Hall effect or, as we show in this Letter, naturally appear in altermagnets due to momentum-dependent spin splitting, see Eq. (5). The nontrivial spatial distribution of the electric potential (spin density) acts as a source of the spin (electric) current vorticity. Therefore, by applying a voltage difference to the contacts, a circulating spin current is expected. In what follows, we confirm this via numerical calculations using two complementary techniques for the semiclassical and fully quantum regimes, respectively.
Figure 2: Electric current $j_{\rm el}=j_{\uparrow}+j_{\downarrow}$ distribution without altermagnetism, $t_{1}=0$ and $t_{2}=0$ (panel (a)). There are no qualitative changes in the shape of the electric current streamlines at nonzero $t_{1}$ or $t_{2}$. Spin current $j_{\rm sp}=j_{\uparrow}-j_{\downarrow}$ distribution in altermagnets for $t_{1}=t_{0}/2$ and $t_{2}=0$ (panel (b)), $t_{1}=0$ and $t_{2}=t_{0}/2$ (panel (c)), and $t_{1}=t_{2}=t_{0}/(2\sqrt{2})$ (panel (d)); configuration of altermagnetic Fermi surfaces is depicted below each of the panels. In all panels, we fix the electric potentials at the contacts as $e\bar{\phi}_{1,\lambda}/t_{0}=-1$ and $e\bar{\phi}_{2,\lambda}/t_{0}=1$. The currents are normalized by $j_{0}=2e\tau t_{0}\,\mbox{max}{\left\{\left|\rho_{\lambda}^{(0)}\left(\bar{\phi}_{2,\lambda}-\bar{\phi}_{1,\lambda}\right)\right|\right\}}/L$, where the maximal value is taken with respect to the spin projections, and we fixed $\rho^{(0)}_{\uparrow}=\rho^{(0)}_{\downarrow}=\rho^{(0)}/2$. Red lines denote the position of the source and drain.
Transport in finite $d$-wave altermagnets.— Let us evaluate electric and spin response in a $d$-wave altermagnet with $J(\mathbf{k})=t_{1}(k_{x}^{2}-k_{y}^{2})+2t_{2}k_{x}k_{y}$. Then, Eq. (6) acquires a particularly simple form:
$t_{0}\nabla^{2}\bar{\phi}_{\lambda}+\lambda t_{1}\left(\nabla_{x}^{2}-\nabla_{y}^{2}\right)\bar{\phi}_{\lambda}+2\lambda t_{2}\nabla_{x}\nabla_{y}\bar{\phi}_{\lambda}=0.$ (9)
This differential equation should be supplemented with boundary conditions. We fix $\bar{\phi}_{\lambda}$ at the contacts, corresponding to an electric or spin voltage bias,  [^3] A different configuration of the contacts where the source and drain are located at the same surface is considered in the Supplemental Material [39]:
$\bar{\phi}_{\lambda}\left(x=0,y\right)$ $=\bar{\phi}_{1,\lambda}(y),\,\,|y|<w,$ (10)
$\bar{\phi}_{\lambda}\left(x=L_{x},y\right)$ $=\bar{\phi}_{2,\lambda}(y),\,\,|y|<w,$
where $L_{x}$ is the length of the sample in the $x$-direction and $w$ is the width of the contacts located at $x=0$ and $x=L_{x}$ sides of the rectangular sample. The component of the electric and spin current normal to the surface vanishes everywhere outside the contacts. In our numerical calculations, we use the finite element method implemented in Wolfram Mathematica [42].
The electric and spin current densities for a few values of parameters $t_{1}$ and $t_{2}$ at $L_{x}=L_{y}=L$ are shown in Fig. 2 for the case of an applied electric voltage bias. The electric current streamlines are only weakly affected by the altermagnetic parameters, hence, we present only a single plot in Fig. 2(a), which shows a typical distribution of the electric current for the Ohmic transport regime. The spin current, however, is nontrivial and depends strongly on the altermagnetic parameters $t_{1}$ and $t_{2}$, which quantify the magnitude of the spin splitting and the orientation of the altermagnetic nodal planes, see Figs. 2(b)–2(d). First of all, vortices of the spin current are observed at the sides of the contacts, independent of the orientation of the nodal planes. The details of the spin current flow are sensitive to the orientation of the altermagnetic nodal planes. In the case when at least one of the planes is perpendicular to the surface with the contacts (e.g., $t_{1}=0$ and $t_{2}\neq 0$), the spin current forms a loop that begins and ends at the same contact, see Fig. 2(c). The spin current away from the contacts is directed perpendicular to the electric field, whose field lines coincide with those of the electric current in Fig. 2(a), confirming the spin splitter effect proposed in Ref. [43]. The latter paper, however, does not consider finite geometries and nonuniform electric fields.
In the case under consideration, the equations describing the distribution of the electrochemical potential (6) and (10) are symmetric with respect to the interchange of the spin species. Therefore, similar to the swirling spin currents shown in Figs. 2(b)–2(d), one can obtain swirling electric currents if the spin imbalance is created at the contacts. The corresponding currents have the same shape and magnitude as those in Fig. 2 with $j_{\rm el}\leftrightarrow j_{\rm sp}$. The spin imbalance can be generated via, e.g., the spin Hall effect [44, 45, 46] or via spin-pumping [47].
The possibility of generating swirling electric currents instead of the spin currents opens the door to much easier and more straightforward observation of the proposed effects via stray magnetic fields. A similar approach was successfully used in mapping out the electric current profile in the electron hydrodynamics [48, 49, 50, 51, 52].
We provide the distribution of the current-induced magnetic field in Fig. 3 for three configurations of the altermagnetic splitting that allows for the current vortices; see Appendix for the definition of the field.
Figure 3: The spatial dependence of the induced magnetic field $B_{z}(\mathbf{r},z)$ for $z=0.1\,L$ at $t_{1}=t_{0}/2$ and $t_{2}=0$ (panel (a)), $t_{1}=0$ and $t_{2}=t_{0}/2$ (panel (b)), and $t_{1}=t_{2}=t_{0}/(2\sqrt{2})$ (panel (c)). In all panels, we fixed the potentials at the contacts as $e\bar{\phi}_{1,\lambda}/t_{0}=-\lambda$ and $e\bar{\phi}_{2,\lambda}/t_{0}=\lambda$, i.e., only the spin imbalance is applied, and we fixed $\rho^{(0)}_{\uparrow}=\rho^{(0)}_{\downarrow}=\rho^{(0)}/2$.
While we focused on the particular model of $d$-wave altermagnets, other models should have similar distributions of the electric and spin currents as long as they allow for $T_{\lambda,ij}\neq 0$ [39]. The only difference will be in the values of the coefficients $T_{\lambda,ij}$ entering the conductivity tensor (4).
Keldysh approach.— To show that the observed swirling currents are not an artifact of the low-energy model, we consider a fully quantum mechanical lattice model of a $d$-wave altermagnet and apply the Keldysh formalism, which was recently used to visualize the spin splitter effect in a homogeneous electric field [53]. The details of the formalism are presented in the Appendix.
The results of the lattice calculations in the disordered regime show a good qualitative agreement with the current distribution obtained in the kinetic approach, cf. Figs. 2(c) and 4. While the side vortices are hard to resolve, the central vortices look similar in both approaches. Vortical currents can also be observed in the ballistic transport regime, albeit the corresponding vorticity is not unique to altermagnets [39].
Figure 4: Spin current distribution obtained via the Keldysh formalism presented in the Appendix. The contacts of the width $w=10a$ are at $x=0,L_{x}$ surfaces. We use $60\times 60$ lattice with the following parameters: $t_{m}=0.3t$, $\mu=-2.0t$, and the voltage bias $eV=0.2t$.
Material estimates.— The characteristic values of the key quantities, i.e., magnetic field and current, are
$B_{0}[n\mbox{T}]\!$ $=$ $\!\frac{73.9}{\sqrt{1\!-(t_{1}^{2}+t_{2}^{2})/t_{0}^{2}}}\frac{\mu\left[\mbox{meV}\right]\!\tau[\mbox{ps}]\delta\phi[\text{mV}]}{L[\mu\text{m}]},$ (11)
$j_{0}\left[\frac{\mbox{A}}{\mbox{m}}\right]\!$ $=$ $\!\frac{0.06}{\sqrt{1\!-(t_{1}^{2}+t_{2}^{2})/t_{0}^{2}}}\frac{\mu\left[\mbox{meV}\right]\!\tau[\mbox{ps}]\delta\phi[\text{mV}]}{L[\mu\text{m}]},$ (12)
where $\delta\phi=\bar{\phi}_{2}-\bar{\phi}_{1}$ is the applied voltage difference.
To provide an estimate of the observed effects, we use the material parameters of such altermagnetic candidates as RuO2  [^4] While the recent studies suggest that RuO2 is nonmagnetic [62, 63, 64], it may support altermagnetism in thin films due to strain [65, 66]. and KRu4O8. We use $t_{0}a^{-2}=2.5~{}\mbox{eV}$, $t_{1}/t_{0}=t_{2}/t_{0}=0.2$ and $\mu=0.4~{}\mbox{eV}$ for RuO2 [4, 55, 56] as well as $t_{1}/t_{0}=t_{2}/t_{0}=0.36$ and $\mu=0.05~{}\mbox{eV}$ for KRu4O8 [9, 56].
[…]
Summary and conclusion.— In this Letter, we showed that the altermagnetic spin splitting in finite-size samples with point contacts allows for swirling electric and spin currents even in disordered systems, i.e., in the Ohmic transport regime. The shape of the vortices encodes the relative orientation of the sample boundaries and the altermagnetic nodal planes, see Fig. 2. Swirling electric (spin) currents can be generated if the spin imbalance (voltage) is applied to the contacts, as is supported by Eqs. (7) and (8). The nontrivial distribution of the electric current is manifested in the induced magnetic field and is within reach of modern magnetometry techniques.
Comparing semiclassical calculations in a continuum model with the Keldysh approach on a lattice, we see qualitatively similar results supporting the model-independence of our findings, cf. Figs. 2(c) and 4. We emphasize that our results do not require strong electron-electron interaction (hydrodynamic transport regime) or ultrapure samples (ballistic transport regime), and take place in the Ohmic transport regime with a fully local response to external stimuli. This broadens the selection of material platforms.
Acknowledgements.
Acknowledgments. A.A.H. acknowledges support from the Project No. 0122U000887 of the Department of Physics and Astronomy of the NAS of Ukraine. The work of E.V.G. was supported by the Program “Dynamics of particles and collective excitations in high-energy physics, astrophysics and quantum macrosystems” (No. 0121U109612) of the Department of Physics and Astronomy of the NAS of Ukraine. This work was supported by the Research Council of Norway through Grant No. 353894 and its Centres of Excellence funding scheme Grant No. 262633 “QuSpin.” Support from Sigma2 - the National Infrastructure for High Performance Computing and Data Storage in Norway, project NN9577K, is acknowledged. P.O.S. acknowledges useful communications with Alireza Qaiumzadeh and Xanthe Verbeek, as well as is grateful to the QuSpin center of excellence for warm hospitality.