### v2

# Electric and spin current vortices in altermagnets

Abstract
Altermagnets constitute a class of collinear magnets with momentum-dependent spin splitting and vanishing net magnetization. Direct observation of the characteristic altermagnetic spin splitting, however, remains challenging. Indirect signatures can be obtained via transport studies, which so far have only considered homogeneous driving fields. We propose to leverage nonuniform electric fields and spin density gradients to probe the shape and the spin polarization of altermagnetic Fermi surfaces via transport measurements. By using both a semiclassical Boltzmann approach and a lattice Keldysh formalism, we show that altermagnets excite swirling electric and spin currents whose profiles depend on the relative orientation of altermagnetic lobes with respect to the sample boundaries. These currents can be measured via magnetometry techniques. Unlike previous proposals considering the hydrodynamic regime of transport, swirling currents are observed even in the Ohmic regime and rely exclusively on the altermagnetic spin splitting, with no swirls observed in ferromagnets. The electric and spin current vortices predicted here provide a different altermagnetic signature in an experimentally accessible setup.
Introduction. Transport of charge and spin represents one of the fundamental types of measurements in condensed matter physics that is indispensable in identifying and investigating new materials. Among such materials that have recently attracted much attention are magnetic materials with anisotropic nonrelativistically spin-split energy bands [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Known as altermagnets [2, 8], such materials are characterized by a combined symmetry including lattice rotations and spin reversal. Altermagnets preserve the inversion symmetry but break the time-reversal symmetry (TRS).
Altermagnetism was predicted in several materials such as MnTe, a few atom layers of RuO2, Mn5Si3, MnF2, V2Se2O, V2Te2O, CrO, CrSb, and intercalated transition metal dichalcogenides CoNb4Se8, and $\kappa$-type organic antiferromagnets [8, 11, 12]. Signatures of spin-split energy bands were observed in angle-resolved photoemission spectroscopy (ARPES) [13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and circular dichroism [23]. However, ARPES measurements generically suffer from low resolution and strong reliance on the comparison between the experimental data and the ab initio calculations.
Transport experiments in altermagnets are mostly focused on the observation of the anomalous Hall effect (AHE) [24, 25, 26]; see also Refs. [27, 28] for earlier discussions of the AHE in altermagnetic candidates. The AHE, however, may originate from different sources, such as a reduced symmetry at the interface [29], domain walls [30], or spin canting [31, 32], see also Ref. [33]. Therefore, the AHE alone cannot be used to unambiguously prove the altermagnetic nature of a material. The spin-splitter effect [34, 35, 36], which is the generation of the spin current by an applied electric field, suggests a different alternative to the AHE for altermagnets. However, its observed signatures are limited to spin-splitter torque in RuO2 [36], whose magnetic state is still debated. Thus, other types of probes are beneficial to rigorously prove the realization of the altermagnetic state.
In this Letter, we propose to use point contacts in altermagnets to generate swirling electric and spin currents whose patterns are directly linked to the symmetry of the altermagnetic spin splitting and the geometry of the sample. While vortices in electron fluids are usually considered as a key indicator of electron hydrodynamics [37], swirling currents in our work appear in the Ohmic transport regime. The proposed effects can be readily investigated via available experimental methods. In particular, the predicted swirling electric currents can be probed via quantum spin magnetometry [38, 39], whereas the distribution of the electrochemical potential can be measured via scanning tunneling potentiometry [40]. Local spin currents can also be probed via techniques similar to the non-local spin valve measurement [41, 42]. Thus, our results complement previous transport measurements and provide an opportunity to cross-verify electric and spin currents via a different set of techniques.
Continuum model and kinetic equations. We start with defining the low-energy continuum model of altermagnets. We focus on two-dimensional (2D) altermagnets as they provide the most direct way to visualize electric and spin currents; the generalization of our formalism to three-dimensional cases is straightforward.
We employ the following model of an altermagnet with a momentum-dependent Zeeman field  [^1] As we verify in Sec. S1.B.2 (see Supplemental Material [46]), a model with sublattice degrees of freedom leads to the same structure of the response tensor.:
$H=t_{0}k^{2}+\sigma_{z}J(\mathbf{k})-\hat{\mu}$. (1)
Here, $\mathbf{k}$ is the momentum vector, $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$, $\sigma_{z}$ is the Pauli matrix in the spin space, $J(\mathbf{k})$ is an effective exchange field, $\hat{\mu}=\text{diag}\left(\mu_{+},\mu_{-}\right)$, where $\mu_{\lambda}$ is the spin-resolved chemical potential for the spin projection $\lambda=\pm$  [^2] The spin-dependent effective chemical potential can be achieved by applying a weak magnetic field via the Zeeman term. In 2D systems, orbital effects can be ignored if the field is in the plane of the material., and parameter $t_{0}$ corresponds to the inverse effective mass of quasiparticles. To obtain finite Fermi surfaces, we assume that $t_{0}k^{2}>\left|J(\mathbf{k})\right|$.
The dispersion relation of electron quasiparticles is
$\varepsilon_{\lambda}=t_{0}k^{2}+\lambda J(\mathbf{k}).$ (2)
The corresponding Fermi surfaces for a $d$-wave altermagnet are schematically shown by the red ($\lambda=+$) and blue ($\lambda=$-) ellipses in Fig. 1.
Figure 1: Schematic setup of a magnet with point-like contacts. The streamlines illustrate the flow of electric (spin) currents induced by the applied spin density (electric potential) difference at the contacts. The upper half corresponds to a $d$-wave altermagnet and the lower half represents a ferromagnet; the corresponding Fermi surfaces with spin-up and spin-down polarized bands are shown in red and blue. Despite the system being diffusive and non-interacting, current vortices arise in the altermagnetic case, unlike the ferromagnetic one.
To address the transport properties of the continuum model, we employ the standard semiclassical approach in which the dynamics of quasiparticles is determined by the Boltzmann kinetic equation,
$\partial_{t}f_{\lambda}+\left(\mathbf{v}_{\mathbf{k},\lambda}\cdot\bm{\nabla}f_{\lambda}\right)+\left(e\mathbf{E}\cdot\partial_{\mathbf{k}}f_{\lambda}\right)=I_{\text{col}}\left\{f_{\lambda}\right\}.$ (3)
Here $f_{\lambda}=f_{\lambda}(t,\mathbf{r},\mathbf{k})$ is the distribution function of quasiparticles wit the spin $\lambda$, $\mathbf{v}_{\mathbf{k},\lambda}=\partial_{\mathbf{k}}\varepsilon_{\mathbf{\lambda}}$ is the quasiparticle velocity, $e<0$ is the electron charge, $\mathbf{E}$ is an electric field, and $I_{\text{col}}\left\{f_{\lambda}\right\}$ is the collision integral which describes scattering off disorder.
For simplicity, we assume nonmagnetic disorder and use the standard relaxation time approximation $I_{\text{col}}\left\{f_{\lambda}\right\}=-\left(f_{\lambda}-\langle f_{\lambda}\rangle\right)/\tau$, where $\tau$ is the relaxation time and $\langle f_{\lambda}\rangle$ is the averaged over the Fermi-surface distribution function. Since we assumed nonmagnetic elastic impurity scattering, the spin-flip processes are determined by the spin-orbital coupling (see, e.g., Ref. [45]) and, therefore, are expected to be weak in altermagnets. Weak spin-flip processes reduce the magnitude of the current but do not qualitatively change the current distribution (see the Supplementary Material [46]).
The charge and current densities in a 2D system for quasiparticles with the spin projection $\lambda$ are
$\left\{\rho_{\lambda}(t,\mathbf{r}),\mathbf{j}_{\lambda}(t,\mathbf{r})\right\}=e\int\frac{d^{2}k}{(2\pi)^{2}}\left\{1,\mathbf{v}_{\mathbf{k},\lambda}\right\}f_{\lambda}(t,\mathbf{r},\mathbf{k}).$ (4)
Electric and spin currents are obtained as $\mathbf{j}_{\rm el}(t,\mathbf{r})=\sum_{\lambda}\mathbf{j}_{\lambda}(t,\mathbf{r})$ and $\mathbf{j}_{\rm sp}(t,\mathbf{r})=\sum_{\lambda}\lambda\,\mathbf{j}_{\lambda}(t,\mathbf{r})$.
By integrating the Boltzmann equation (3) over momenta, we derive the continuity equation
$\partial_{t}\rho_{\lambda}+\left(\bm{\nabla}\cdot\mathbf{j}_{\lambda}\right)=0.$ (5)
Multiplying Eq. (3) by $e\mathbf{v}_{\mathbf{k},\lambda}$ and integrating over momenta, we obtain
$\partial_{t}j_{\lambda,j}+\nabla_{i}\tilde{\Pi}_{ij}+e\Pi_{ij}E_{i}=-\frac{j_{\lambda,j}}{\tau},$ (6)
where
$\tilde{\Pi}_{ij}(t,\mathbf{r})$ $=$ $e\int\frac{d^{2}k}{(2\pi)^{2}}v_{\mathbf{k},\lambda,i}v_{\mathbf{k},\lambda,j}f_{\lambda}(t,\mathbf{r},\mathbf{k}),$ (7)
$\Pi_{ij}(t,\mathbf{r})$ $=$ $-e\int\frac{d^{2}k}{(2\pi)^{2}}(\partial_{k_{i}}\partial_{k_{j}}\varepsilon_{\lambda})f_{\lambda}(t,\mathbf{r},\mathbf{k}).$ (8)
Assuming weak deviations from equilibrium, neglecting temperature deviations, and linearizing Eq. (6), we derive the following equation for the current density:
$\partial_{t}j_{\lambda,l}+\frac{\sigma_{il}}{\tau}\nabla_{i}\bar{\phi}_{\lambda}=-\frac{j_{\lambda,l}}{\tau}.$ (9)
In writing this equation, we introduced the conductivity tensor
$\sigma_{ij}=2e\tau\rho^{(0)}_{\lambda}\left(t_{0}\delta_{ij}+\lambda T_{\lambda,ij}\right),$ (10)
where
$T_{\lambda,ij}=\frac{e}{2\rho^{(0)}_{\lambda}}\int\frac{d^{2}k}{(2\pi)^{2}}\left[\partial_{k_{i}}\partial_{k_{j}}J(\mathbf{k})\right]f_{\lambda}^{(0)}(\mathbf{k})$ (11)
is the part of the conductivity tensor determined by the effective exchange field, $f^{(0)}_{\lambda}(\mathbf{k})=1/\left[e^{\left(\varepsilon_{\lambda}-\mu_{\lambda}\right)/T}+1\right]$ is the equilibrium Fermi-Dirac distribution function, $\rho^{(0)}_{\lambda}$ is the equilibrium charge density, and $e\bar{\phi}_{\lambda}=\delta\mu_{\lambda}+e\phi$ is the effective spin-resolved electrochemical potential with $\phi$ being electric potential and $\delta\mu_{\lambda}$ being the spin-resolved deviation of the chemical potential from its equilibrium value.
Equation (9) is general and can be applied both for static and dynamic perturbations. In this Letter, we focus on the stationary transport regime, where Eq. (9) reduces to
$\nabla_{j}\left(\sigma_{ij}\nabla_{i}\bar{\phi}_{\lambda}\right)=0.$ (12)
This differential equation should be supplemented with boundary conditions. For example, one can fix $\bar{\phi}_{\lambda}$ at the contacts or require a constant injected current. Then, Eq. (12) is a partial differential equation that can be solved numerically or, in certain cases, such as a ribbon with the fixed values of current at the point-like contacts (see the Supplemental Material (SM) [46]), analytically.
Vorticity in the Ohmic transport regime. Before delving into the numerical solutions to Eq. (12), let us present intuitive arguments supporting the possibility of vortices in the Ohmic transport regime. We start with the analog of the vorticity equation in fluid mechanics [47] and consider the spin-resolved analog of vorticity $\bm{\omega}_{\lambda}=\bm{\nabla}\times\mathbf{j}_{\lambda}$. Assuming the absence of the equilibrium spin density, $\mu_{\lambda}=\mu$, the equations for the electric $\bm{\omega}_{\rm el}=\sum_{\lambda}\bm{\omega}_{\lambda}$ and spin $\bm{\omega}_{\rm sp}=\sum_{\lambda}\lambda\bm{\omega}_{\lambda}$ vorticities follow from Eq. (9)
$\partial_{t}\omega_{{\rm el},j}+\frac{\omega_{{\rm el},j}}{\tau}=-2e\rho^{(0)}\epsilon_{jlm}T_{im}\nabla_{l}\nabla_{i}\bar{\phi}_{\rm sp},$ (13)
$\partial_{t}\omega_{{\rm sp},j}+\frac{\omega_{{\rm sp},j}}{\tau}=-2e\rho^{(0)}\epsilon_{jlm}T_{im}\nabla_{l}\nabla_{i}\bar{\phi}_{\rm el},$ (14)
where $\epsilon_{jlm}$ is the fully antisymmetric tensor. These equations allow us to identify two sources of vorticity: (i) off-diagonal elements of the altermagnetic part of the conductivity tensor $T_{\lambda,i\neq m}$ and (ii) anisotropy in the diagonal elements $T_{\lambda,ii}\neq T_{\lambda,mm}$ with $i\neq m$. The former allows for the vorticity even when potentials change only along one direction, e.g., for rectangular contacts. The off-diagonal terms can originate, e.g., from the Hall effect or, as we show in this Letter, naturally appear in altermagnets due to momentum-dependent spin splitting, see Eq. (11). Vorticity for different diagonal elements, which also appear in altermagnets, requires nonuniform potential with $\nabla_{l}\nabla_{i}\bar{\phi}_{\lambda}\neq 0$ at $l\neq i$. The latter occurs at the edges of rectangular contacts or for point-like contacts. In both cases, nontrivial spatial distribution of the electric potential (spin density) acts as a source of the spin (electric) current vorticity. Therefore, by applying a voltage difference to the contacts, a circulating spin current is expected. In what follows, we confirm this via numerical calculations using two complementary techniques for the semiclassical and fully quantum regimes, respectively.
Figure 2: (a) Electric current $j_{\rm el}=j_{\uparrow}+j_{\downarrow}$ distribution without altermagnetism, $t_{1}=0$ and $t_{2}=0$. There are no qualitative changes in the shape of the electric current streamlines at nonzero $t_{1}$ or $t_{2}$. Spin current $j_{\rm sp}=j_{\uparrow}-j_{\downarrow}$ distribution in altermagnets for (b) $t_{1}=t_{0}/2$ and $t_{2}=0$, (c) $t_{1}=0$ and $t_{2}=t_{0}/2$, and (d) $t_{1}=t_{2}=t_{0}/(2\sqrt{2})$; the configuration of altermagnetic Fermi surfaces is depicted below each of the panels. In all panels, we fix the electric potentials at the contacts as $e\bar{\phi}_{1,\lambda}/t_{0}=-1$ and $e\bar{\phi}_{2,\lambda}/t_{0}=1$. The currents are normalized by $j_{0}=2e\tau t_{0}\,\mbox{max}{\left\{\left|\rho_{\lambda}^{(0)}\left(\bar{\phi}_{2,\lambda}-\bar{\phi}_{1,\lambda}\right)\right|\right\}}/L$, where the maximal value is taken with respect to the spin projections, and we fixed $\rho^{(0)}_{\uparrow}=\rho^{(0)}_{\downarrow}=\rho^{(0)}/2$. Red lines denote the position of the source and drain.
Transport in finite $d$-wave altermagnets. Let us evaluate electric and spin response in a $d$-wave altermagnet with $J(\mathbf{k})=t_{1}(k_{x}^{2}-k_{y}^{2})+2t_{2}k_{x}k_{y}$. Then, Eq. (12) acquires a particularly simple form,
$t_{0}\Delta\bar{\phi}_{\lambda}+\lambda t_{1}\left(\nabla_{x}^{2}-\nabla_{y}^{2}\right)\bar{\phi}_{\lambda}+2\lambda t_{2}\nabla_{x}\nabla_{y}\bar{\phi}_{\lambda}=0.$ (15)
This differential equation should be supplemented with boundary conditions. We fix $\bar{\phi}_{\lambda}$ at the contacts, corresponding to an electric or spin voltage bias  [^3] A different configuration of the contacts where the source and drain are located at the same surface is considered in the Supplemental Material [46],
$\bar{\phi}_{\lambda}\left(x=0,y\right)$ $=\bar{\phi}_{1,\lambda}(y),\,\,|y|<w,$ (16)
$\bar{\phi}_{\lambda}\left(x=L_{x},y\right)$ $=\bar{\phi}_{2,\lambda}(y),\,\,|y|<w,$
where $L_{x}$ is the length of the sample in the $x$-direction and $w$ is the width of the contacts located at $x=0$ and $x=L_{x}$ sides of the rectangular sample. The component of the electric and spin current normal to the surface vanishes everywhere outside the contacts. In our numerical calculations, we use the finite element method implemented in Wolfram Mathematica [49].
The electric and spin current densities for a few values of parameters $t_{1}$ and $t_{2}$ at $L_{x}=L_{y}=L$ are shown in Fig. 2 for the case of an applied electric voltage bias. The electric current streamlines are only weakly affected by the altermagnetic parameters, and hence we present only a single plot in Fig. 2(a), which shows a typical distribution of the electric current for the Ohmic transport regime. The spin current, however, is nontrivial and depends strongly on the altermagnetic parameters $t_{1}$ and $t_{2}$, which quantify the magnitude of the spin splitting and the orientation of the altermagnetic nodal planes, see Figs. 2(b)–2(d). First of all, vortices of the spin current are observed at the sides of the contacts, independent of the orientation of the nodal planes. These vortices originate from the anisotropy of $T_{\lambda,ii}$. The details of the spin current flow are sensitive to the orientation of the altermagnetic nodal planes. In the case when at least one of the planes is perpendicular to the surface with the contacts (e.g., $t_{1}=0$ and $t_{2}\neq 0$), the spin current forms a loop that begins and ends at the same contact, see Fig. 2(c). This type of swirling current relies on the off-diagonal components of $T_{\lambda,ij}$. The spin current away from the contacts is directed perpendicular to the electric field, whose field lines coincide with those of the electric current in Fig. 2(a), confirming the spin splitter effect [34, 35]. These papers, however, do not consider finite geometries and nonuniform electric fields.
In the case under consideration, the equations describing the distribution of the electrochemical potential (12) and (16) are symmetric with respect to the interchange of the spin species. Therefore, similar to the swirling spin currents shown in Figs. 2(b)–2(d), one can obtain swirling electric currents if the spin imbalance is created at the contacts. The corresponding currents have the same shape and magnitude as those in Fig. 2 with $j_{\rm el}\leftrightarrow j_{\rm sp}$. The spin imbalance can be generated via, e.g., the spin Hall effect [50, 51, 52] or via spin-pumping [53].
The possibility of generating swirling electric currents instead of the spin currents opens the door to the observation of the proposed effects via stray magnetic fields. A similar approach was successfully used in mapping out the electric current profile in the electron hydrodynamics [54, 55, 56, 57, 58]. We provide the distribution of the current-induced magnetic field in Fig. 3 for three configurations of the altermagnetic splitting that allows for the current vortices; see the SM [46] for the definition of the field.
Figure 3: The spatial dependence of the induced magnetic field $B_{z}(\mathbf{r},z)$ for z=0.1Lz=0.1\,L at (a) $t_{1}=t_{0}/2$ and $t_{2}=0$, (b) $t_{1}=0$ and $t_{2}=t_{0}/2$, and (c) $t_{1}=t_{2}=t_{0}/(2\sqrt{2})$. In all panels, we fixed the potentials at the contacts as $e\bar{\phi}_{1,\lambda}/t_{0}=-\lambda$ and $e\bar{\phi}_{2,\lambda}/t_{0}=\lambda$, i.e., only the spin imbalance is applied, and we fixed $\rho^{(0)}_{\uparrow}=\rho^{(0)}_{\downarrow}=\rho^{(0)}/2$.
While we focused on the particular model of $d$-wave altermagnets, other models of altermagnets should have similar distributions of the electric and spin currents as long as they allow for off-diagonal or anisotropic diagonal components of the altermagnetic part of the conductivity tensor (see the SM [46]). The only difference will be in the values of the coefficients $T_{\lambda,ij}$ entering the conductivity tensor (10).
Keldysh approach. To show that the observed swirling currents are not an artifact of the low-energy model, we consider a fully quantum mechanical lattice model of a $d$-wave altermagnet and apply the Keldysh formalism [59, 60], which was recently used to visualize the spin splitter effect in a homogeneous electric field [61].
A $d_{xy}$-altermagnet is modeled by the following tight-binding Hamiltonian:
$H_{S}=\sum_{i\sigma}\varepsilon_{i}c^{\dagger}_{i\sigma}c_{i\sigma}+\sum_{ij\sigma\sigma^{\prime}}c_{i\sigma}^{\dagger}t_{ij}^{\sigma\sigma^{\prime}}c_{j\sigma^{\prime}},$ (17)
where $c_{i\sigma}$ ($c_{i\sigma}^{{\dagger}}$) is the annihilation (creation) operator of electrons with the spin $\sigma$ at the site $i$. The onsite potential is $\varepsilon_{i}$ and the spin-dependent hopping terms
$\hat{t}_{ij}=\begin{cases}-t\mathbb{I},\quad\left(j=i\pm e_{x}\;\mathrm{or}\;j=i\pm e_{y}\right)\\ -t_{m}\sigma_{z},\quad\left[j=i\pm\left(e_{x}+e_{y}\right)\right]\\ t_{m}\sigma_{z},\quad\left[j=i\pm\left(e_{x}-e_{y}\right)\right],\end{cases}$ (18)
where $e_{i}$ is the unit vector in the $i$-th direction and $t_{m}$ corresponds to the strength of the spin-dependent diagonal hopping. The leads and connection between the leads and sample region are modelled as metals with $\varepsilon_{i}=0$, and we only consider nearest neighbor hopping $t_{ij}^{\mathrm{leads}}=-t\mathbb{I}$ and $t_{ij}^{\mathrm{interface}}=-t\mathbb{I}$, respectively. The constrictions between the leads and the sample are modelled by a very high on-site potential $\varepsilon_{\mathrm{edge}}=10^{5}\,t$ on the edges of the sample, except for a region of $10$ sites where the on-site potential is zero.
We define a bond spin-current operator representing the flow of spin $S_{k}$ from site $i$ to $j$ via
$J_{ij}^{S_{k}}\equiv\frac{1}{4i}\sum_{\alpha\beta}\left(c_{j\beta}^{\dagger}\{\sigma_{k},\hat{t}_{ji}\}_{\beta\alpha}c_{i\alpha}-\mathrm{H.c.}\right),$ (19)
where the curly brackets denote the anticommutator. The steady-state non-equilibrium statistical average of this current is calculated in the Keldysh formalism (see the SM [46]).
To include disorder in the altermagnet, we uniformly assign the fraction $n_{I}=0.2$ of the sites with a local potential of the strength $\varepsilon_{\rm imp}=10\,t_{m}$ simulating pointlike impurities. The disorder average is taken with respect to $100$ configurations of impurities.
The results of the lattice calculations in the disordered regime show a good qualitative agreement with the current distribution obtained in the kinetic approach; cf. Figs. 2(c) and 4. While the side vortices are hard to resolve due to the low magnitude of the current there, the central vortices look similar in both approaches. Vortical currents can also be observed in the ballistic transport regime, albeit the corresponding vorticity is not unique to altermagnets (see the SM [46]).
Figure 4: Spin current distribution obtained via the Keldysh formalism. The contacts of the width w=10aw=10\,a are at $x=0,L_{x}$ surfaces. We use $60\times 60$ lattice with the following parameters: $t_{m}=0.3\,t$, $\mu=-2.0\,t$, and the voltage bias eV=0.2teV=0.2\,t.
Material estimates. The characteristic values of the key quantities, i.e., magnetic field and current, are
$B_{0}[n\mbox{T}]\!$ $=$ $\!\frac{73.9}{\sqrt{1\!-(t_{1}^{2}+t_{2}^{2})/t_{0}^{2}}}\frac{\mu\left[\mbox{meV}\right]\!\tau[\mbox{ps}]\delta\phi[\text{mV}]}{L[\mu\text{m}]},$ (20)
$j_{0}\left[\frac{\mbox{A}}{\mbox{m}}\right]\!$ $=$ $\!\frac{0.06}{\sqrt{1\!-(t_{1}^{2}+t_{2}^{2})/t_{0}^{2}}}\frac{\mu\left[\mbox{meV}\right]\!\tau[\mbox{ps}]\delta\phi[\text{mV}]}{L[\mu\text{m}]},$ (21)
where $\delta\phi=\bar{\phi}_{2}-\bar{\phi}_{1}$ is the applied voltage difference.
To estimate the observed effects, we use the material parameters of altermagnetic candidates such as RuO2  [^4] While the recent studies suggest that RuO2 is nonmagnetic [76, 77, 78], it may support altermagnetism in thin films due to strain [79, 19]. and KRu4O8. We use $t_{0}a^{-2}=2.5~\mbox{eV}$, $t_{1}/t_{0}=t_{2}/t_{0}=0.2$, and $\mu=0.4~\mbox{eV}$ for RuO2 [4, 63, 64] and $t_{1}/t_{0}=t_{2}/t_{0}=0.36$ and $\mu=0.05~\mbox{eV}$ for KRu4O8 [9, 64].
Assuming the width of ribbon $L=5~\mu\mbox{m}$, typical scattering time $\tau=1~\mbox{ps}$, and fixing the applied potential difference $\delta\phi=1~\mbox{mV}$, we estimate $B_{0}\approx 6~\mu\mbox{T}$ and the current density $j_{0}\approx 24\,\mu\mbox{A/m}$ for RuO2 as well as $B_{0}\approx 1~\mu\mbox{T}$ and the current density $j_{0}=3\,\mu\mbox{A/m}$ for KRu4O8. These values are small but are within reach of modern magnetic field imaging techniques such as nanowire magnetic force microscopy [65, 66], the scanning superconducting quantum interference device (SQUID) magnetometry [67, 66], and also at the edge of the sensitivity of the quantum spin magnetometry [68, 38, 66, 69]. In addition to the magnetic field, one can measure the local distribution of the electric potential, albeit the latter is less sensitive to the nontrivial distribution of currents [46].
Summary and conclusion. In this Letter, we showed that the altermagnetic spin splitting in finite-size samples with point contacts allows for swirling electric and spin currents even in disordered systems; i.e., in the Ohmic transport regime. The shape of the vortices encodes the relative orientation of the sample boundaries and the altermagnetic nodal planes, see Fig. 2. Swirling electric (spin) currents can be generated if the spin imbalance (voltage) is applied to the contacts, as is supported by Eqs. (13) and (14). The nontrivial distribution of the electric current is manifested in the induced magnetic field and is within reach of modern magnetometry, complementing previous transport measurements.
Comparing semiclassical calculations in a continuum model with the Keldysh approach on a lattice, we see qualitatively similar results supporting the model-independence of our findings; cf. Figs. 2(c) and 4. We emphasize that our results do not require strong electron-electron interaction (hydrodynamic transport regime) or ultrapure samples (ballistic transport regime), and take place in the Ohmic transport regime with a fully local response to external stimuli. This broadens the selection of material platforms.
Acknowledgements.
Acknowledgments. A.H. acknowledges support from Project No. 0122U000887 of the Department of Physics and Astronomy of the NAS of Ukraine. The work of E.V.G. was supported by the Program “Dynamics of particles and collective excitations in high-energy physics, astrophysics and quantum macrosystems” (Grant No. 0121U109612) of the Department of Physics and Astronomy of the NAS of Ukraine. This work was supported by the Research Council of Norway through Grant No. 353894 and its Centres of Excellence funding scheme Grant No. 262633 “QuSpin.” Support from Sigma2 - The National Infrastructure for High Performance Computing and Data Storage in Norway, Project NN9577K, is acknowledged. P.O.S. acknowledges useful communications with Alireza Qaiumzadeh and Xanthe Verbeek, and is grateful to the QuSpin center of excellence for warm hospitality.